The inter-atomic spacing of a linear 1D lattice of atoms of mass 8.0 ×10^-27 kg is
5.0 Å. The force constant is 3.0×10^-2 N/m
. Calculate the maximum frequency
which can be supported by this lattice
1
Expert's answer
2020-02-04T09:22:54-0500

The maximum frequency can be supported by such 1D lattice is


"\omega_\text{max}=\sqrt{\frac{4C}{M}}=\sqrt{\frac{4\cdot3\cdot10^{-2}}{8.0 ×10^{-27}}}=3.87\text{ rad/s}."